# The steep redshift evolution of the hierarchical binary black hole merger rate may cause the $z$-$\chi_{\rm eff}$ correlation

### v2

@@ Appendix D Mass ratio distributions
Our inferred mass ratio distributions are influenced both by our inferred primary and secondary component mass distributions, and by our inferred pairing function. The peak of our pairing function is inferred to be at $\mu^{q}=0.87^{+0.07}_{-.07}$. This results in the mass ratio distributions shown in Fig. 7. The subpopulation’s mass ratio distribution peaks at lower values than that of the main population, consistent with expectations for 2G+1G mergers.
However, the value at which it peaks depends strongly on our choice of pairing function. Assuming the same pairing function between the main and subpopulation – as we do in Fig. 7 and all other figures in the main text – results in a mass ratio distribution that peaks at $q=0.73$. Assuming a different pairing function between the main and subpopulations results in a mass ratio distribution that peaks at $\approx 0.5$, which is in line with conventional expectations from 2G+1G mergers (Rodriguez et al., 2019; Ye et al., 2025). However, the data are not yet able to constrain the pairing function of the subpopulation, so this result is likely spurious.
@@ Appendix F Additional figures
Fig. 12 displays the 99th percentiles of the component mass distributions in Fig. 2. It would seem to imply that the pair-instability mass feature begins at $\approx 46\,\mathrm{M}_{\odot}$. This is consistent with Wang et al. (2022), Pierra et al. (2024), and Antonini et al. (2025), but appears inconsistent with Wang et al. (2025), who find a population of low-spin objects extending to $\approx 70\,\mathrm{M}_{\odot}$. To directly compare to this finding, we calculate the fraction of systems in the spinning subpopulation in various mass bins and find that this fraction lower in the range $50\,\mathrm{M}_{\odot},70\,\mathrm{M}_{\odot}]$ than in the range $70\,\mathrm{M}_{\odot},90\,\mathrm{M}_{\odot}]$ to 91.56%. This supports the hypothesis put forth in Wang et al. (2025), but at low significance.

### v3

@@ Appendix D Mass ratio distributions
Our inferred mass ratio distributions are influenced both by our inferred primary and secondary component mass distributions, and by our inferred pairing function. The peak of our pairing function is inferred to be at $\mu^{q}=0.85^{+0.07}_{-.06}$. This results in the mass ratio distributions shown in Fig. 7. The subpopulation’s mass ratio distribution peaks at lower values than that of the main population, consistent with expectations for 2G+1G mergers.
However, the value at which it peaks depends strongly on our choice of pairing function. Assuming the same pairing function between the main and subpopulation – as we do in Fig. 7 and all other figures in the main text – results in a mass ratio distribution for the subpopulation that peaks at $q=0.73$. Assuming a different pairing function between the main and subpopulations results in a mass ratio distribution that peaks at $\approx 0.5$, which is in line with conventional expectations from 2G+1G mergers (Rodriguez et al., 2019; Ye et al., 2025). However, the data are not yet able to constrain the pairing function of the subpopulation, so this result may be spurious.
@@ Appendix F Additional figures
Fig. 12 displays the 99th percentiles of the component mass distributions in Fig. 2. It would seem to imply that the pair-instability mass feature begins at $\approx 46\,\mathrm{M}_{\odot}$. This is consistent with Wang et al. (2022), Pierra et al. (2024), and Antonini et al. (2025), but appears inconsistent with Wang et al. (2025), who find a population of low-spin objects extending to $\approx 70\,\mathrm{M}_{\odot}$. To directly compare to this finding, we calculate the fraction of systems in the spinning subpopulation in various mass bins and find that this fraction is lower in the range $m_{1}\in[50\,\mathrm{M}_{\odot},70\,\mathrm{M}_{\odot}]$ than in the range $m_{1}\in[70\,\mathrm{M}_{\odot},90\,\mathrm{M}_{\odot}]$ to 91.1%. This supports the hypothesis put forth in Wang et al. (2025), but at low significance.